OpenStudy (anonymous):

A common inhabitant of human intestines is Ecoli. The cell divides every twenty minutes and the initial population is 60. Find the rate of growth after 8 hours

OpenStudy (paxpolaris):

do you want it in cells per minute or cells per hour?

OpenStudy (anonymous):

per hour please

OpenStudy (paxpolaris):

Can you get the equation of the Population P(t) at any given time t

OpenStudy (anonymous):

P(t)= 60* e^(k*t) where k=2.0794

OpenStudy (paxpolaris):

how did you get 2.07?

OpenStudy (paxpolaris):

nevermind it's correct

OpenStudy (paxpolaris):

so you then need to find the equation for P'(t) and solve for P'(8)

OpenStudy (anonymous):

how? @PaxPolaris

OpenStudy (paxpolaris):

\[P(t)= 60e^{kt}...\ where\ k=\ln(2^3)\]\[P'(t)=60ke^{kt}\]